# DRAGON: Distributional Rewards Optimize Diffusion Generative Models

### v1

@@ Appendix D Comparison with Open-Source Models
  • MusicGen (Copet et al., 2023), a non-diffusion auto-regressive music generation model that predicts discrete audio tokens (Défossez et al., 2024) in sequence, coming in small, medium, and large variants.
  • Stable Audio Open (Evans et al., 2024a), a state-of-the-art audio diffusion model that generates variable durations of up to 45 seconds, following a similar design to our base model.
@@
We use the non-vocal subset of Song Describer (SDNV) as the evaluation benchmark, and compare in aesthetics score, CLAP score (with text), per-song FAD, and dataset FAD metrics. Based on the results shown in Figure 10, we make the following observations:
@@ Appendix E Model Details and Hyperparameters
Convolutional VAE. We build on the Descript Audio Codec (DAC, or Improved RVQGAN) (Kumar et al., 2023) architecture and training scheme by using a KL-bottleneck with a dimension of 32 and an effective hop of 768 samples, resulting in an approximately 57Hz VAE. We train to convergence using the recommended mel-reconstruction loss and the least-squares GAN formulation with $\ell_{1}$ feature matching on multi-period and multi-band discriminators.
@@
Diffusion Transformer (DiT). Following the base model in (Novack et al., 2025), our model backbone builds upon DiT-XL (Peebles & Xie, 2023), with modifications aimed at optimizing computational efficiency. Specifically, it uses a streamlined transformer block design, consisting of a single attention layer followed by a single feed-forward layer, similar to Llama (Dubey et al., 2024). The diffusion hyperparameter design follows EDM (Karras et al., 2022), with $\sigma_{\textrm{data}}=0.5$, $P_{\textrm{mean}}=-0.4$, $P_{\textrm{std}}=1.0$, $\sigma_{\textrm{max}}=80$, and $\sigma_{\textrm{min}}=0.002$. Also following EDM, we apply a logarithmic transformation to the noise levels, followed by sinusoidal embeddings. These processed noise-level embeddings are then combined and integrated into the DiT block through an adaptive layer normalization block. For text conditioning, we concatenate the T5-embedded text tokens with audio tokens at each attention layer. As a result, the audio token query attends to a concatenated sequence of audio and text keys, enabling the model to jointly extract relevant information from both modalities. Pre-training of the baseline diffusion model lasted five days across 32 Nvidia A100 GPUs with a total batch size of 256 and a learning rate of $10^{-4}$ with Adam.
@@
DRAGON training settings and hyperparameters. All DRAGON fine-tuning is performed on top of the baseline model introduced above on four or eight A100 GPUs with a total batch size of 80, with the baseline model used as the “reference model $f_{\textrm{ref}}$” required by the loss functions, as defined in the $A_{\theta}$ term in eq. 7. During DRAGON fine-tuning, the textual conditions are from the pre-training dataset, and no ground-truth audio is used unless required by the reward function. We use Adam with a fixed learning rate of $3\times 10^{-6}$ and a gradient clip of $45$ (determined via gradient logging). For the DPO loss eq. 6 and the KTO loss eq. 8, we select $\beta=5000$ following (Wallace et al., 2024) and do not update $f_{\textrm{ref}}$. With paired DPO or KTO, the batch of $80$ demonstrations is generated from 40 prompts in pairs, each pair consisting of one positive and one negative demonstration. With unpaired KTO, the 80 demonstrations are from 80 distinct prompts, and the positive/negative label is assigned by comparing each demonstration with the batch mean. All training runs (pre-training and DRAGON) use a 10% condition dropout to enhance classifier-free guidance (CFG). The online audio demonstrations in DRAGON are generated with the default inference setting (40 diffusion steps with the second-order DPM sampler with CFG++ enabled in selected time steps). We use $f_{\theta}$ to produce the demonstrations with probability $0.9$ and use $f_{\textrm{ref}}$ with probability $0.1$. Intuitively, mixing in $f_{\textrm{ref}}$ generations “anchors” the generation quality to the reference model and provides additional regularization.
@@ 16 D_pos, D_neg = compare_r_dist(
For reward functions that evaluate distributions, we show a piece of psuedocode with GPU parallelization logic. The function compare_r_dist handles the parallelization and calls optimize_r_dist which implements Algorithm 1.
@@ 24 D_pos, D_neg = optimize_r_dist(
For Algorithm 1 (optimize_r_dist), we show an example implementation that optimizes full-dataset FAD as follows. Other rewards like Vendi can be handled similarly.
@@ Appendix H Full Result Tables
We use five tables to list the reward metrics achieved by all DRAGON models discussed in the paper. First, we present the win rates and the average values of instance-level rewards (aesthetics score, CLAP score, and per-song FAD) in Tables 6 and 7. Next, we present the win rates and the average values of distribution-level rewards like $r_{\textrm{dist}}$ (full-dataset FAD and Vendi), evaluated in a bootstrapped setting, in Tables 8 and 9. As mentioned in Section 5.1, this means sampling 40-example generation subsets from our 2185-prompt evaluation set with replacement 1000 times, computing the reward metric for each subset, and reporting the average and win rate among these 1000 numbers. Finally, we present the distribution-level rewards evaluated over the full 2185-instance evaluation set without bootstrapping in Table 10. Vendi is computed over per-song average MA embeddings in the full-dataset setting as in the main paper body, but is computed over all MA embeddings without averaging in the bootstrapped setting.

### v2

@@ Appendix D Comparison with Open-Source Models
  • MusicGen (MusicGen), a non-diffusion auto-regressive music generation model that predicts discrete audio tokens (defossez2023high) in sequence, coming in small, medium, and large variants.
  • Stable Audio Open (StableAudio), a state-of-the-art audio diffusion model that generates variable durations of up to 45 seconds, following a similar design to our base model.
@@
We use the non-vocal subset of Song Describer (SDNV) as the evaluation benchmark, and compare in aesthetics score, CLAP score (with text), per-song FAD, and dataset FAD metrics. Based on the results shown in Figure˜10, we make the following observations:
@@ Appendix E Model Details and Hyperparameters
Convolutional VAE. We build on the Descript Audio Codec (DAC, or Improved RVQGAN) (DAC) architecture and training scheme by using a KL-bottleneck with a dimension of 32 and an effective hop of 768 samples, resulting in an approximately 57Hz VAE. We train to convergence using the recommended mel-reconstruction loss and the least-squares GAN formulation with $\ell_{1}$ feature matching on multi-period and multi-band discriminators.
@@
Diffusion Transformer (DiT). Following the base model in (Presto), our model backbone builds upon DiT-XL (DiT), with modifications aimed at optimizing computational efficiency. Specifically, it uses a streamlined transformer block design, consisting of a single attention layer followed by a single feed-forward layer, similar to Llama (Llama3). The diffusion hyperparameter design follows EDM (EDM), with $\sigma_{\textrm{data}}=0.5$, $P_{\textrm{mean}}=-0.4$, $P_{\textrm{std}}=1.0$, $\sigma_{\textrm{max}}=80$, and $\sigma_{\textrm{min}}=0.002$. Also following EDM, we apply a logarithmic transformation to the noise levels, followed by sinusoidal embeddings. These processed noise-level embeddings are then combined and integrated into the DiT block through an adaptive layer normalization block. For text conditioning, we concatenate the T5-embedded text tokens with audio tokens at each attention layer. As a result, the audio token query attends to a concatenated sequence of audio and text keys, enabling the model to jointly extract relevant information from both modalities. Pre-training of the baseline diffusion model lasted five days across 32 Nvidia A100 GPUs with a total batch size of 256 and a learning rate of $10^{-4}$ with Adam.
@@
DRAGON training settings and hyperparameters. All DRAGON fine-tuning is performed on top of the baseline model introduced above on four or eight A100 GPUs with a total batch size of 80, with the baseline model used as the “reference model $f_{\textrm{ref}}$” required by the loss functions, as defined in the $A_{\theta}$ term in eq.˜7. Fine-tuning takes 3-5 days depending on the reward, but we found this can be accelerated by reducing the number of diffusion steps. During DRAGON fine-tuning, the textual conditions are from the pre-training dataset, and no ground-truth audio is used unless required by the reward function. We use Adam with a fixed learning rate of $3\times 10^{-6}$ and a gradient clip of $45$ (determined via gradient logging). For the DPO loss eq.˜6 and the KTO loss eq.˜8, we select $\beta=5000$ following (Diffusion-DPO) and do not update $f_{\textrm{ref}}$. With paired DPO or KTO, the batch of $80$ demonstrations is generated from 40 prompts in pairs, each pair consisting of one positive and one negative demonstration. With unpaired KTO, the 80 demonstrations are from 80 distinct prompts, and the positive/negative label is assigned by comparing each demonstration with the batch mean. All training runs (pre-training and DRAGON) use a 10% condition dropout to enhance classifier-free guidance (CFG). The online audio demonstrations in DRAGON are generated with the default inference setting (40 diffusion steps with the second-order DPM sampler with CFG++ enabled in selected time steps). We use $f_{\theta}$ to produce the demonstrations with probability $0.9$ and use $f_{\textrm{ref}}$ with probability $0.1$. Intuitively, mixing in $f_{\textrm{ref}}$ generations “anchors” the generation quality to the reference model and provides additional regularization. We use this same configuration for each reward studied, selecting set sizes to fill GPU memory.
@@ 16 D_pos, D_neg = compare_r_dist(
For reward functions that evaluate distributions, we show a piece of psuedocode with GPU parallelization logic. The function compare_r_dist handles the parallelization and calls optimize_r_dist which implements Algorithm˜1.
@@ 24 D_pos, D_neg = optimize_r_dist(
For Algorithm˜1 (optimize_r_dist), we show an example implementation that optimizes full-dataset FAD as follows. Other rewards like Vendi can be handled similarly.
@@ Appendix H Full Result Tables
We use five tables to list the reward metrics achieved by all DRAGON models discussed in the paper. First, we present the win rates and the average values of instance-level rewards (aesthetics score, CLAP score, and per-song FAD) in Tables 7 and 8. Next, we present the win rates and the average values of distribution-level rewards like $r_{\textrm{dist}}$ (full-dataset FAD and Vendi), evaluated in a bootstrapped setting, in Tables 9 and 10. As mentioned in Section˜5.1, this means sampling 40-example generation subsets from our 2185-prompt evaluation set with replacement 1000 times, computing the reward metric for each subset, and reporting the average and win rate among these 1000 numbers. Finally, we present the distribution-level rewards evaluated over the full 2185-instance evaluation set without bootstrapping in Table˜11. Vendi is computed over per-song average MA embeddings in the full-dataset setting as in the main paper body, but is computed over all MA embeddings without averaging in the bootstrapped setting.